OpenStudy (anonymous):

help will reward medals The radioactive substance strontium-90 has a half-life of 28 years. In other words, it takes 28 years for half of a given quantity of strontium-90 to decay to a non-radioactive substance. The amount of radioactive strontium-90 still present after t years is modeled by the expression grams. Evaluate the expression for t = 0, t = 28 and t = 56. What does each value of the expression represent?

OpenStudy (anonymous):

expression is 80*2^-t/28

OpenStudy (anonymous):

@eliassaab @michelleg46 @lgbasallote @shadowfiend and others help

OpenStudy (anonymous):

@shadowfiend @anonymoustwo44 @lgbasallote @rebeccaskell94 @razor99 help

OpenStudy (anonymous):

@dumbcow @dpaInc

OpenStudy (phi):

As you know \(2^{1} =2 \) and \(2^{2} = 2\cdot 2\) More tricky is \( 2^{-1}= \frac{1}{2^1} =\frac{1}{2} \) and \(2^{-2} = \frac{1}{2\cdot 2}\) (A negative exponent means "flip upside down" and make the exponent positive. A very important fact is \(2^0=1 \) Any number (except 0) to the zero power equals 1